OpenStudy (anonymous):

you toss 4 coins in the air at the same time what is the probability of each of the classes occurring 4 heads 3heads: 1 tail 2heads: 2 tails 1head: 3 tails 4tails

OpenStudy (anonymous):

even if though there are multiple combinations for the i figured the 4 heads and the 4 tails had a 1/16 chance of happening but i screwed up somewhere with the middle three because my over all chance of occurring comes out to 1.5. sorry i should have explained better earlier. 3heads: 1 tail 2heads: 2 tails 1head: 3 tails

OpenStudy (dan815):

1 3 6 3 1 = 14 total ways and 4 heads = 1/14 4 tails = 1/14 3 heads 1 tail = 3/14 3 tails 1 head = 3/14 2 heads 2 tails = 6/14

OpenStudy (anonymous):

ok thanks that explains where i messed up

OpenStudy (dan815):

there is a mistake 1 4 6 4 1 = 16 total ways and 4 heads = 1/16 4 tails = 1/16 3 heads 1 tail = 4/16 3 tails 1 head = 4/16 2 heads 2 tails = 6/16

OpenStudy (dan815):

there is anotehr way

OpenStudy (dan815):

- - - - 4 spots you will fill up with heads and tails which more than 1/2 success rate so (1/2)^2 how every a different number of combinations that result in same outcome for 3 heads 1 tail it is 4 choose 1 for 2 heads 2 tails it is 4Choose for 4 heads, and 4 tilts its 4choose 0

OpenStudy (dan815):

soo 1/16*4choose0 = prob of 4 heads or 4 tails

OpenStudy (dan815):

1/16*4choose1 = prob of (3,1) 1/6 * 4C2 = prob of (2,2)

OpenStudy (anonymous):

Awesome bro

OpenStudy (anonymous):

@dan815